Weather - overcast most of the day, warm and humid, but a strong southerly breeze helped keep the humidity and bugs in check.
John went with Gordon in the morning to change batteries at the 4 southwest sites. Sounding launched at 1800 UTC without issue and reached 59.4 hPa before bursting. John spent the afternoon doing testing on the SODAR amplifier modules as outlined by Eckhard. Confirmed that modules C1 and D1 were not functional, there are no short circuits in the speaker array, and that the amplifiers were being keyed/biased only during the transmit pulse. Module D1 was damaged during testing. No root cause of the excessive heat on the amplifier mounting rails that was noticed lasted week was found. John will replace the two missing amplifier plate mounting screws just for good measure, i.e. to help with heat sinking. Removed Voyager disc CHEESEHEAD_1 and inserted CHEESEHEAD_2 into drive. John will bring CHEESEHEAD_1 back to Boulder on Sunday.
Finally, a medium to large garter snake was found in the SODAR trailer, near the cable port in the front left of the trailer, behind the UPS that is on the floor. When startled with a stick, it exited the trailer thru the cable port in the floor, but will likely return. We are looking for some steel wool to plug the hole with in the trailers. If anyone knows where some is, please contact the senior ISS technician.